# Topological characterization of magnon-polaron bands and thermal Hall conductivity in a frustrated kagome antiferromagnet

### v1

@@ A.1 Decoupling magnon-phonon modes by canonical transformation
are independent of the position in the lattice, the terms $X_{q}^{l(nl)}$ ($q\in 1,2,3,4$) appearing in the exponents depend explicitly on the positions of the magnons. For clarity and conciseness, we present the explicit forms of $X_{q}^{l(nl)}$ for a representative bond within a single unit cell. To demonstrate this, we consider $a_{\tilde{r}-\tilde{\delta}_{3}}^{\dagger}b_{\tilde{r}}e^{X_{1}^{l(nl)}}$ and $a_{\tilde{r}-\tilde{\delta}_{3}}^{\dagger}b_{\tilde{r}}^{\dagger}e^{X_{2}^{l(nl)}}$ as representative NN interactions, and $a_{\tilde{r}-\tilde{\delta}_{3}}^{\dagger}b_{\tilde{r}-2\tilde{\delta}_{2}}e^{X_{3}^{l(nl)}}$ and $a_{\tilde{r}-\tilde{\delta}_{3}}^{\dagger}b_{\tilde{r}-2\tilde{\delta}_{2}}^{\dagger}e^{X_{4}^{l(nl)}}$ as NNN interactions. For the local interaction, the corresponding trems $X_{q}^{l}$ can be written as,
@@
Although the terms $X_{q}^{l(nl)}$ differ for other NN or NNN bonds, the number of phonon degrees of freedom remains the same in each of the cases. Throughout the calculations presented in our work, we have neglected the four-magnon quartic terms such as $m_{i}^{\dagger}m_{j}^{\dagger}m_{i}m_{j}$.
@@ A.2 Reduced Hamiltonian via finite-temperature phonon averaging
Now, to eliminate the phonon degrees of freedom, we take a finite-temperature phonon average of the Hamiltonian in Eq. (22), namely,
$\tilde{H}_{R}=\braket{\tilde{H}}_{\mathrm{ph}}.$ (26)
Here, $\tilde{H}$ contains both phonon operators and magnon operators, and only the phonon operators take part in this averaging. To further elaborate on the method, we consider a generic phonon operator of the form,
$\widetilde{O}=e^{\gamma(\tilde{b}_{r}^{\dagger}-\tilde{b}_{r})},$ (27)
where $\tilde{b}_{r}$ is the phonon operator corresponds to the lattice site $r$, $\gamma$ denotes the coupling strength, and the associated phonon energy of that site is $H_{\mathrm{ph}}=\hbar\omega(\tilde{b}_{r}^{\dagger}\tilde{b}_{r})$. Then, the phonon average can be expressed as
@@
where |$\ket{n}$ is the $n$-th state of harmonic oscillator, $\beta=1/k_{B}T$ and $f_{\mathrm{ph}}=1/(e^{\beta\hbar\omega}-1)$ is the BE distribution function. Now the expectation of $\widetilde{O}$ in Eq. (28) yields [103, 104]
@@
$\braket{\widetilde{O}}=\left[e^{\beta\hbar\omega}f_{\mathrm{ph}}\right]^{-1}e^{-\frac{\gamma^{2}}{2}}\sum_{n=0}^{\infty}e^{-n\beta\hbar\omega}\mathcal{L}_{n}(\gamma^{2})=e^{-\frac{1}{2}\gamma^{2}(2f_{\mathrm{ph}}+1)},$ (31)
where $\mathcal{L}_{n}(\gamma^{2})$ is the Laguerre polynomial function. Further, applying the thermal averaging in Eq. (26), the reduced Hamiltonian assumes the form,
$\tilde{H}_{R}=S$ $\left[\sum_{i}\left(4(J_{1}+J_{2})\,{\mathcal{E}}+4D\,{\mathcal{E}}_{1}+\mu_{0}B_{0}\sin{\eta}-\Delta^{l(nl)}\right)m_{i}^{\dagger}m_{i}\right.$
@@
$\lambda_{1}^{l}$ $=4\alpha^{2}(2f_{\text{ph}}+1),~{}\lambda_{2}^{l}=8\alpha^{2}(2f_{\text{ph}}+1),~{}\lambda_{3}^{l}=\lambda_{4}^{l}=6\alpha^{2}(2f_{\text{ph}}+1),$
$\lambda_{1}^{nl}$ $=\lambda_{3}^{nl}=5\alpha^{2}(2f_{\text{ph}}+1),\lambda_{2}^{nl}=\lambda_{4}^{nl}=3\alpha^{2}(2f_{\text{ph}}+1).$ (33)
@@ Appendix B Phase diagram for a higher value of $J_{2}$
In the main text, we focused on results obtained for small values of $J_{2}$ and $D$, since a finite $J_{2}(>0)$ along with a moderate DMI can stabilize the $120^{\circ}$ coplanar spin configuration in KAFM. Additionally, these parameters play a crucial role in determining the emergence of non-coplanar ground state ordering in the presence of an external magnetic field. Hence, to provide a more comprehensive view of the parameter space, we present the corresponding phase diagrams at a larger NNN exchange interaction strength, $J_{2}=0.1J_{1}$, as functions of $D$ and $g^{l}$ in Fig.11(a), and $D$ and $g^{nl}$ in Fig.11(b).
Figure 11: The combined phase diagrams are plotted for $J_{2}=0.1J_{1}$, $B_{0}=0.4B_{s}$, $\widetilde{T}=0$, as a function of (a) $D$ and $g^{l}$, and (b) $D$ and $g^{nl}$.
@@ Appendix C Band Structures at finite temperature
Figure 12: In the presence of local spin-phonon coupling, the band structures are shown for (a) $g^{l}=0.8J_{1},~{}\widetilde{T}=0.3$ and (b) $g^{l}=0.8J_{1},~{}\widetilde{T}=0.7$. Further, in the presence of non-local spin-phonon coupling, the band structures are shown for (c) $g^{nl}=0.8J_{1},~{}\widetilde{T}=0.3$ and (d) $g^{nl}=0.8J_{1},~{}\widetilde{T}=0.7$. The other parameters are fixed at $B_{0}=0.4B_{s}$, $J_{2}=0.03J_{1}$, $D=0.045J_{1}$, $\hbar\omega=J_{1}S$.
In presence of local spin-phonon coupling, we observe that the two lower bands (that is lower and middle) exhibit a positive energy shift and the band gap between them increases near the high symmetry points (see Figs. 12(a) and 12(b)) as a function of temperature. In contrast, in the presence of the non-local coupling, the band gap near the $\mathbf{K}$ point decreases, which raises the possibility of another gap-closing transition at higher temperatures. However, here also the two lower bands exhibit a positive energy shift as shown in Figs. 12(c) and 12(d). Thus, if magnon-polaron instability arises at higher values of the phonon frequency $\omega$, the situation can be stabilized via raising the temperature. However, since magnon fluctuations, decay processes, and scattering become significant at high temperatures and have been neglected in this study, the stabilization is effective only in the low-temperature regime.

### v2

@@ A.1 Decoupling magnon-phonon modes by canonical transformation
are independent of the position in the lattice, the terms $X_{q}^{l(nl)}$ ($q\in 1,2,3,4$) appearing in the exponents depend explicitly on the positions of the magnons. For clarity and conciseness, we present the explicit forms of $X_{q}^{l(nl)}$ for a representative bond within a single unit cell. To demonstrate this, we consider $a_{\tilde{r}-\tilde{\delta}_{3}}^{\dagger}b_{\tilde{r}}e^{X_{1}^{l(nl)}}$ and $a_{\tilde{r}-\tilde{\delta}_{3}}^{\dagger}b_{\tilde{r}}^{\dagger}e^{X_{2}^{l(nl)}}$ as representative NN interactions, and $a_{\tilde{r}-\tilde{\delta}_{3}}^{\dagger}b_{\tilde{r}-2\tilde{\delta}_{2}}e^{X_{3}^{l(nl)}}$ and $a_{\tilde{r}-\tilde{\delta}_{3}}^{\dagger}b_{\tilde{r}-2\tilde{\delta}_{2}}^{\dagger}e^{X_{4}^{l(nl)}}$ as NNN interactions. For the local interaction, the corresponding terms $X_{q}^{l}$ can be written as
@@
Although the terms $X_{q}^{l(nl)}$ differ for other NN or NNN bonds, the number of phonon degrees of freedom remains the same in each of the cases. Throughout the calculations presented in our work, we have neglected the four-magnon quartic terms such as $m_{i}^{\dagger}m_{j}^{\dagger}m_{i}m_{j}$ ($\forall i,j$).
@@ A.2 Reduced Hamiltonian via finite-temperature phonon averaging
Now, to eliminate the phonon degrees of freedom, we take a finite-temperature phonon average of the transformed Hamiltonian (Eq. (23)), namely
$\widetilde{H}_{\text{R}}=\braket{\widetilde{H}}_{\mathrm{ph}}.$ (27)
Here, $\widetilde{H}$ contains both phonon operators and magnon operators, and only the phonon operators take part in this averaging. To further elaborate on the method, we consider a generic phonon operator of the form
$\widetilde{O}=e^{\alpha(\tilde{b}_{r}^{\dagger}-\tilde{b}_{r})},$ (28)
where $\tilde{b}_{r}$ is the phonon operator corresponds to the lattice site $r$, $\alpha$ denotes the coupling strength, and the associated phonon energy of that site is $H_{\mathrm{ph}}=\hbar\omega(\tilde{b}_{r}^{\dagger}\tilde{b}_{r})$. Then, the thermal phonon average of $\widetilde{O}$ with respect to the $n$-th harmonic oscillator state, namely |$\ket{n}$ can be expressed as
@@
where $\beta=1/k_{B}T$ and $f_{\mathrm{ph}}=1/(e^{\beta\hbar\omega}-1)$ is the BE distribution function. Now, the expectation of $\widetilde{O}$ in Eq. (29) yields [104, 105]
@@
$\braket{\widetilde{O}}=\left[e^{\beta\hbar\omega}f_{\mathrm{ph}}\right]^{-1}e^{-\frac{\alpha^{2}}{2}}\sum_{n=0}^{\infty}e^{-n\beta\hbar\omega}\mathcal{L}_{n}(\alpha^{2})=e^{-\frac{1}{2}\alpha^{2}(2f_{\mathrm{ph}}+1)},$ (32)
where $\mathcal{L}_{n}(\alpha^{2})$ is the Laguerre polynomial function. Further, applying the thermal averaging in Eq. (27), the reduced Hamiltonian is obtained as
$\widetilde{H}_{\text{R}}=S$ $\left[\sum_{i}\left(4(J_{1}+J_{2})\,{\mathcal{E}}+4D\,{\mathcal{E}}_{1}+\mu_{B}B_{0}\sin{\eta}-\Delta^{l(nl)}\right)m_{i}^{\dagger}m_{i}\right.$
@@
$\lambda_{1}^{l}$ $=4(\alpha^{l})^{2}(2f_{\text{ph}}+1),~{}\lambda_{2}^{l}=8(\alpha^{l})^{2}(2f_{\text{ph}}+1),~{}\lambda_{3}^{l}=\lambda_{4}^{l}=6(\alpha^{l})^{2}(2f_{\text{ph}}+1),$
$\lambda_{1}^{nl}$ $=\lambda_{3}^{nl}=5(\alpha^{nl})^{2}(2f_{\text{ph}}+1),\lambda_{2}^{nl}=\lambda_{4}^{nl}=3(\alpha^{nl})^{2}(2f_{\text{ph}}+1).$ (34)
@@ Appendix B Phase diagram for a higher value of $J_{2}$
In the main text, we focused on results obtained for small values of $J_{2}$ and $D$, since a finite $J_{2}(>0)$ along with a moderate out-of-plane DMI can stabilize the $120^{\circ}$ coplanar spin configuration in KAFM. Additionally, these parameters play a crucial role in determining the emergence of non-coplanar ground state ordering in the presence of an external magnetic field.
Hence, to provide a more comprehensive view of the parameter space, we present the corresponding phase diagrams at a larger coupling strength, $J_{2}=0.1J_{1}$, as functions of $D$ and $g^{l}$ in Fig.12(a), and $D$ and $g^{nl}$ in Fig.12(b). Although in both cases, the phase plots emerge with the same number of distinct topological phases, the phase boundaries differ from each other owing to different coupling mechanisms. Notably, for these higher values of $J_{2}$ and $D$, the corresponding $\omega_{c}$ would be higher than the values shown in Fig. 2. Hence, we keep $\hbar\omega=J_{1}S$ for both the cases.
Figure 12: The combined phase diagrams are presented for $J_{2}=0.1J_{1}$, $B_{0}=0.4B_{s}$, $\widetilde{T}=0$, $\hbar\omega=J_{1}S$ in the (a) $g^{l}$-$D$ and (b) $g^{nl}$-$D$ plane, where both $g^{l(nl)}$ and $D$ are measured in units of $J_{1}$. the Chern number set ($C_{1},C_{2},C_{3}$) is designated for each of the topological phases.
@@ Appendix C Band Structures at finite temperature
In the presence of local spin-phonon coupling, we observe that the two lower bands (that is lower and middle) exhibit a positive energy shift and the band gap between them increases near the high symmetry points (see Figs. 13(a) and 13(b)) as a function of temperature. In contrast, in the presence of the non-local coupling, the band gap near the $\mathbf{K}$ point decreases, which raises the possibility of another gap closing transition at higher temperatures. However, here also the two lower bands exhibit a positive energy shift as shown in Figs. 13(c) and 13(d). Thus, if magnon-polaron instability arises at higher values of the phonon frequency $\omega$, the situation can be stabilized via raising the temperature.
However, since magnon fluctuations, decay processes, and scattering become significant at high temperatures ($k_{B}T\gtrsim J_{1}S$), which have been neglected in this study, the stabilization is effective only in the low-temperature regime.
Figure 13: In the presence of local spin-phonon coupling, the band structures (with the corresponding Chern numbers highlighted in yellow) are shown for (a) $g^{l}=0.8J_{1},~{}\widetilde{T}=0.3$ and (b) $g^{l}=0.8J_{1},~{}\widetilde{T}=0.7$. Further, in the presence of non-local spin-phonon coupling, the band structures are shown for (c) $g^{nl}=0.8J_{1},~{}\widetilde{T}=0.3$ and (d) $g^{nl}=0.8J_{1},~{}\widetilde{T}=0.7$. The other parameters are fixed at $B_{0}=0.4B_{s}$, $J_{2}=0.03J_{1}$, $D=0.045J_{1}$ and $\hbar\omega=J_{1}S$.